A 30% acid solution is to be mixed with a 50% acid solution to get 100 milliliters of a 35% acid solution. How many milliliters of 30% are needed
Answer by amoresroy(361)

A 30% acid solution is to be mixed with a 50% acid solution to get 100 milliliters of a 35% acid solution. How many milliliters of 30% are needed
Let x = the number of milliliters of 30% needed
100-x = the number of milliliters of 50% needed
0.30x + 0.50(100-x) = .35(100)
0.30x + 50 - 0.5x = 35
-0.2x = 35-50
x = 75

75 milliliters of 30% are needed